1. Consider a perfectly competitive market where the market demand curve is given by Q 72-4P and the market supply curve is gPlease answer question B
Answer #1

A tax of 3 will reduce the demand and so new demand equation is Q = 72 - 4(P + 3) or Q = 60 - 4P. Supply equation is Q = -6 + 2P. New equilibrium price and quantity are

60 - 4P = -6 + 2P

66 = 6P

P(sellers) = 66/6 = $11

P(buyers) = 11 + 3 = $14

Quantity = -6 + 2*11 = 16 units